# Edge Learning for 6G-enabled Internet of Things: A Comprehensive Survey of Vulnerabilities, Datasets, and Defenses

### v1

@@ IV-B2 In-Application Domains
- The URLLC Target is 1 millisecond round-trip latency
@@
- Supports large number of devices (up to 1 million per square kilometer), low data rates (tens of kbps),
@@
- Dealing with potential interference issues with large scale devices.
@@
IV-C 5G/6G peer-to-peer IoT communications
With the advent of next-generation networking technologies such as 5G and the forthcoming 6G, we’re about to witness an unparalleled transformation in IoT communications. These new-generation networks promise exponential improvements in speed, capacity, and latency, which are vital for efficient and reliable IoT functionalities. This sub-section provides an exhaustive review of the emerging technologies in 5G and projected 6G networking that is poised to revolutionize peer-to-peer IoT communications. Table IX presents the new technologies in 5G/6G IoT communications.
IV-D 5G IoT Technologies
@@
  • Ultra-Reliable Low-Latency Communications (URLLC): URLLC is a crucial part of 5G technology. It aims to provide ultra-reliable communication links with low latency, which is important for critical applications such as autonomous vehicles, remote surgeries, and industrial automation [103].
@@
  • Network Slicing: Network slicing is a 5G feature that allows the creation of multiple virtual networks over a common physical infrastructure. This allows for better resource allocation and quality of service management, especially important in IoT applications where different devices may have vastly different requirements [105].
@@
IV-E 6G IoT Technologies
@@ IV-F1 Infrastructural Measures
Integrating training and inference functionalities across multiple nodes in the network helps to maintain privacy and confidentiality, and ensure a high level of security. Considering that the 6G infrastructure is designed for connected intelligence and will deploy AI at different network hierarchy levels, the AI-powered mini-cell can be used to block certain threats such as DoS, man-in-the-middle, and information theft at the lowest levels before it reaches the targeted applications [5].
@@
To preserve privacy, Over-the-Air FL is getting attention as a prominent approach for fast wireless data aggregation through exploiting the property of overlay in multiple access channels, without the need for accessing the network nodes’ private data [86]. The context-aware security is a promising area of research where the goal is to obtain a network that can extract the threat level from the existing situation, use the context to determine the necessary security level, and match security levels to security parameters [88].
@@
Another intelligent security concept that is predicted to play an important role in future networks is the Moving Target Defense (MTD). MTD technologies can provide enhanced security flexibility by continuously and dynamically reshaping the underlying system on different layers during the execution cycle. This makes it difficult for adversaries to successfully exploit a continuously evolving, instantaneous, and unpredictable targeted system. For example, when an attack is noticed, the MTD module(s) can tell the SDN controller which way to re-program its data plane to reduce the extent of the attack [44]. Furthermore, various security and privacy architectural mechanisms are proposed [92], including the protection of sensitive customer information such as location, for which privacy-aware ML-based offloading schemes can be used. In addition, biometrics-based authentication or a password-less utility can be adopted within the network’s access control arrangements.
IV-F2 Additional Services
@@
Apart from the native security intelligence expected for 6G, including the techniques mentioned above, Nguyen et al. [44] envisioned that the big improvement in 6G security is the broad adoption of security-targeted AI-driven techniques and in-line deep packet inspection (DPI) capabilities within firewalls and network intrusion detection systems (NIDS). As an example, the authors suggested that the 5G security gateway situated on the Access and Mobility Function (AMF) will require significant enhancements to its existing capabilities which they predict, among other things, in 6G will involve an integrated AI-based intrusion prevention engine. Therefore, other advanced technologies and services predicted to protect the security and privacy of future 6G mobile networks have virtually secured their positions. For example, Blockchain and Distributed Ledger Technologies (DLT) are expected to support next-generation networks, including 6G, in terms of privacy, transparency, and immutability in various critical tasks. These tasks include access control, authentication, mutual trust, anonymization, and single point of failure perturbations avoidance at the infrastructure and application levels [111, 2].
@@
Current Blockchain designs and operational smartness security implementations are mainly realized in two forms, including, 1) programmable software and 2) consensus mechanism. The programmable software is called a smart contract, which automatically performs operations according to predefined conditions. The consensus mechanism, which is basically a distributed algorithm that verifies the validity of submitted candidate blocks, with the verification process being different depending on the type of Blockchain (i.e., public, private, or consortium) as well as its implementation logic (e.g. puzzle-based, vote-based, etc.).
@@
The 6G is expected to introduce ultra-high speeds (over 1Tbps [112]), large data volumes (5 zettabytes per month by 2030 [113]), and a huge number of globally connected mobile devices (13.1 billion by 2023 [114]). AI-based security should be implemented in these operations to keep up with the 6G, and to enable the identification of anomalous patterns in blockchain transactions and potential vulnerabilities in smart contracts at higher speeds and low costs [44].
IV-G Highlights 5G/6G in Peer-to-Peer IoT Learning
@@ IV-G1 Network Slicing
IV-G2 Increased Speed and Capacity
@@
IV-G3 Ultra-Low Latency
Another distinct feature of 5G/6G is its ultra-low latency, which can be reduced to as low as 1 millisecond. This latency minimization is essential for IoT devices, permitting them to have rapid peer-to-peer learning [99].
IV-G4 Reliability and Availability
@@
IV-G5 Edge Computing Support
5G has significantly enhanced edge computing, optimizing cloud computing systems by processing data close to the network’s edge, where data generation occurs. This local data processing capability means that IoT devices can learn from data on-site, reducing the need for long-distance data transmission and hastening peer-to-peer learning. Edge Intelligence uses AI methods, which are anticipated to be a crucial factor in the development of future 6G networks, bolstering their performance, enabling new services, and introducing new functions [119].
IV-G6 Energy minimization
Le et al. [120] apply reconfigurable intelligent surface (RIS)-aided wireless power transfer to improve battery life in federated learning (FL)-based wireless networks. Therefore, optimizing transmission time, power control, and the RIS’s phase shifts, shows that the total transmit power can be minimized while meeting both minimum harvested energy and transmission data rate requirements. Zarandi and Tabassum [121] present a federated deep reinforcement learning (DRL) framework using double deep Q-network (DDQN) to optimize multi-objective problems in IoT devices, minimizing task completion delay and energy consumption. The framework enhances scalability and privacy, and the results show faster learning speeds than federated DQN and non-federated DDQN. Zheng et al. [122] proposes an optimization strategy for wireless-powered multi-access edge computing networks to minimize total computation delay. Using a mixed integer programming model, the proposed strategy breaks down the problem into sub-parts for offloading decisions and power transfer duration optimization. In addition, the proposed strategy leverages deep reinforcement learning to handle time-varying channel conditions, which leads to near-minimal computation delays and low computational complexity.
TABLE X: Edge Learning-based Security and Privacy Solutions.

### v2

@@ III-B2 In-Application Domains
- The URLLC Target is 1-millisecond round-trip latency
@@
- Supports the large number of devices (up to 1 million per square kilometer), low data rates (tens of kbps),
@@
- Dealing with potential interference issues with large-scale devices.
@@
III-C 5G/6G peer-to-peer IoT communications
With the advent of next-generation networking technologies such as 5G and the forthcoming 6G, we’re about to witness an unparalleled transformation in IoT communications. These new-generation networks promise exponential speed, capacity, and latency improvements, which are vital for efficient and reliable IoT functionalities. This sub-section provides an exhaustive review of the emerging technologies in 5G and projected 6G networking that is poised to revolutionize peer-to-peer IoT communications. Table VI presents the new technologies in 5G/6G IoT communications.
III-D 5G IoT Technologies
The following five technologies are integral to the functioning of the 5G IoT. We will delve into Enhanced Mobile Broadband (eMBB), which enhances data bandwidth and latency; Ultra-Reliable Low-Latency Communications (URLLC), crucial for critical applications; Massive Machine Type Communications (mMTC), designed to handle large-scale IoT deployments; Network Slicing, a feature that allows for effective resource allocation and quality management; and Edge Computing, a significant factor in reducing latency and network congestion. These technologies collectively revolutionize IoT, facilitating efficient, robust communication and creating various applications, from daily tasks to essential industrial processes.
@@
  • Ultra-Reliable Low-Latency Communications (URLLC): URLLC is a crucial part of 5G technology. It aims to provide ultra-reliable communication links with low latency, which is essential for critical applications such as autonomous vehicles, remote surgeries, and industrial automation [72].
@@
  • Network Slicing: Network slicing is a 5G feature that allows creating multiple virtual networks over a common physical infrastructure. This allows for better resource allocation and quality of service management, especially important in IoT applications where different devices may have vastly different requirements [74].
@@
III-E 6G IoT Technologies
This subsection delves into the future of connectivity with 6G IoT technologies. As we prepare to step into the next generation of networks, we explore the transformative role of Terahertz (THz) Communications that promise unprecedented data rates; Advanced AI and Machine Learning set to manage the network infrastructure and process immense data; Integration of Satellite and Terrestrial Networks offering global coverage; Advanced Positioning and Sensing for precise location tracking and sensor data incorporation; and finally, Advanced Security and Privacy to address escalating concerns surrounding data protection. Together, these technologies encapsulate the advancements that 6G will bring, pushing the boundaries of IoT to new heights.
@@ III-F1 Infrastructural Measures
Integrating training and inference functionalities across multiple network nodes helps maintain privacy and confidentiality and ensures high security. Considering that the 6G infrastructure is designed for connected intelligence and will deploy AI at different network hierarchy levels, the AI-powered mini-cell can be used to block specific threats such as DoS, man-in-the-middle and information theft at the lowest levels before it reaches the targeted applications [8].
@@
To preserve privacy, Over-the-Air FL is getting attention as a prominent approach for fast wireless data aggregation by exploiting the overlay property in multiple access channels, without the need for accessing the network nodes’ private data [55]. The context-aware security is a promising area of research where the goal is to obtain a network that can extract the threat level from the existing situation, use the context to determine the necessary security level, and match security levels to security parameters [57].
@@
Another intelligent security concept predicted to play an essential role in future networks is the Moving Target Defense (MTD). MTD technologies can provide enhanced security flexibility by continuously and dynamically reshaping the underlying system on different layers during the execution cycle. This makes it difficult for adversaries to successfully exploit a continuously evolving, instantaneous, and unpredictable targeted system. For example, when an attack is noticed, the MTD module(s) can tell the SDN controller which way to re-program its data plane to reduce the extent of the attack [58]. Furthermore, various security and privacy architectural mechanisms are proposed [62], including protecting sensitive customer information such as location, for which privacy-aware ML-based offloading schemes can be used. In addition, biometrics-based authentication or a password-less utility can be adopted within the network’s access control arrangements.
III-F2 Additional Services
@@
Apart from the native security intelligence expected for 6G, including the techniques mentioned above, Nguyen et al. [58] envisioned that the significant improvement in 6G security is the broad adoption of security-targeted AI-driven techniques and in-line deep packet inspection (DPI) capabilities within firewalls and network intrusion detection systems (NIDS). For example, the authors suggested that the 5G security gateway situated on the Access and Mobility Function (AMF) will require significant enhancements to its existing capabilities which they predict, among other things, in 6G will involve an integrated AI-based intrusion prevention engine. Therefore, other advanced technologies and services predicted to protect the security and privacy of future 6G mobile networks have virtually secured their positions. For example, Blockchain and Distributed Ledger Technologies (DLT) are expected to support next-generation networks, including 6G, regarding privacy, transparency, and immutability in various critical tasks. These tasks include access control, authentication, mutual trust, anonymization, and single point of failure perturbations avoidance at the infrastructure and application levels [80, 3].
@@
Current Blockchain designs and operational smartness security implementations are mainly realized in two forms, including 1) programmable software and 2) consensus mechanism. The programmable software is an intelligent contract that automatically performs operations according to predefined conditions. The consensus mechanism is a distributed algorithm that verifies the validity of submitted candidate blocks, with the verification process being different depending on the type of Blockchain (i.e., public, private, or consortium) and its implementation logic (e.g., puzzle-based, vote-based, etc.).
@@
The 6G is expected to introduce ultra-high speeds (over 1Tbps [81]), large data volumes (5 zettabytes per month by 2030 [82]), and a vast number of globally connected mobile devices (13.1 billion by 2023 [83]). AI-based security should be implemented in these operations to keep up with the 6G, and to identify anomalous patterns in blockchain transactions and potential vulnerabilities in smart contracts at higher speeds and low costs [58].
III-G Highlights 5G/6G in Peer-to-Peer IoT Learning
@@ III-G1 Network Slicing
III-G2 Increased Speed and Capacity
@@
III-G3 Ultra-Low Latency
Another distinct feature of 5G/6G is its ultra-low latency, which can be reduced to as low as one millisecond. This latency minimization is essential for IoT devices, permitting them to have rapid peer-to-peer learning [68].
III-G4 Reliability and Availability
@@
III-G5 Edge Computing Support
5G has significantly enhanced edge computing, optimizing cloud computing systems by processing data close to the network’s edge, where data generation occurs. This local data processing capability means IoT devices can learn from data on-site, reducing the need for long-distance data transmission and hastening peer-to-peer learning. Edge Intelligence uses AI methods, which are anticipated to be a crucial factor in developing future 6G networks, bolstering their performance, enabling new services, and introducing new functions [88].
III-G6 Energy minimization
Le et al. [89] apply reconfigurable intelligent surface (RIS)-aided wireless power transfer to improve battery life in federated learning (FL)-based wireless networks. Therefore, optimizing transmission time, power control, and the RIS’s phase shifts show that the total transmit power can be minimized while meeting minimum harvested energy and transmission data rate requirements. Zarandi and Tabassum [90] present a federated deep reinforcement learning (DRL) framework using double deep Q-network (DDQN) to optimize multi-objective problems in IoT devices, minimizing task completion delay and energy consumption. The framework enhances scalability and privacy, and the results show faster learning speeds than federated DQN and non-federated DDQN. Zheng et al. [91] proposes an optimization strategy for wireless-powered multi-access edge computing networks to minimize total computation delay. Using a mixed integer programming model, the proposed strategy breaks down the problem into sub-parts for offloading decisions and power transfer duration optimization. In addition, the proposed strategy leverages deep reinforcement learning to handle time-varying channel conditions, which leads to near-minimal computation delays and low computational complexity.
TABLE VII: Edge Learning-based Security and Privacy Solutions.